OpenStudy (anonymous):

To travel 84 miles, it takes Sue, riding a moped, 3 hours less time than it takes Doreen to travel 63 miles riding a bicycle. Sue travels 12 miles per hour faster than Doreen. Find the times and rates of both girls.

OpenStudy (anonymous):

Let s be Sue's speed. Doreen's speed is 12 mph slower than Sue's or s-12. Let t be the time that Doreen travels 63 miles. r t = d Solve the following equations for s and t: {s (t - 3 ) = 84, (s - 12) t = 63} There are two solutions. Use s=21mph and t=7 hours Sue takes 4 hours to travel 84 miles @ 21 mph Doreen takes 7 hours to travel 63 miles @ 9 mph